OpenStudy (anonymous):

If an allele makes up one half of the frequency of a population’s allele for a given trait, its allele frequency is A. 100 percent. B. 75 percent. C. 50 percent. D. 4 percent.

OpenStudy (anonymous):

i have no idea umm B? @uri @M00SE

OpenStudy (uri):

What is Allele..?

OpenStudy (anonymous):

50%

OpenStudy (uri):

@M00SE What is Allele...?

OpenStudy (anonymous):

One way to understand how and why populations evolve is to imagine a model of a hypothetical population that does not evolve. If a popu-lation is not evolving, allele frequencies in its gene pool do not change, which means that the population is in genetic equilibrium. Sexual Reproduction and Allele Frequency Gene shuffling during sexual reproduction produces many gene combinations. But a century ago, researchers realized that meiosis and fertilization, by themselves, do not change allele frequencies. So hypothetically, a population of sexually reproducing organisms could remain in genetic equilibrium. The Hardy- Weinberg Principle The Hardy- Weinberg principle states that allele frequencies in a population should remain constant unless one or more factors cause those frequencies to change. The Hardy- Weinberg principle makes predictions like Punnett squares— but for populations, not individuals. Here’s how it works. Suppose that there are two alleles for a gene: A ( dominant) and a ( recessive).A cross of these alleles can produce three possible genotypes: AA, Aa, and aa. The frequencies of genotypes in the population can be predicted by these equations, where p and q are the frequencies of the dominant and recessive alleles: In symbols: p2 + 2pq + q2 = 1 and p + q = 1 In words: ( frequency of AA) + ( frequency of Aa) + ( frequency of aa) = 100% and ( frequency of A) + ( frequency of a) = 100% Suppose that, in one generation, the frequency of the A allele is 40 per-cent ( p = 0.40) 40per-cent(p=0.40) and the frequency of the a allele is 60 percent ( q = 0.60). LOL?

OpenStudy (anonymous):

In my book (@.@)

OpenStudy (anonymous):

Allele; form of gene: one of two or more alternative forms of a gene, occupying the same position locus on paired chromosomes and controlling the same inherited characteristic

OpenStudy (anonymous):

lol thanks @m00se

OpenStudy (uri):

Won't it be 75? If it is One half?

OpenStudy (anonymous):

It's 50%. Google says so ;-;

OpenStudy (anonymous):

thats wat i tght

OpenStudy (uri):

Let me see,Google sucks.

OpenStudy (anonymous):

grrrr ill let you guys kno

OpenStudy (uri):

Yahoo says 50 too,Idk :3

OpenStudy (anonymous):

ill let u kno lol

OpenStudy (anonymous):

hold on tho it will be a few

OpenStudy (anonymous):

hey @m00se pm me?

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

@m00se was right

OpenStudy (anonymous):

welp @uri was too lol

OpenStudy (uri):

xD

OpenStudy (jtfan2000):

I agree with uri - 75%